Using a =(3,4) and b=(2,-1) as base vectors, express the vector C =(7,2) in form ya+ub



1
Expert's answer
2021-09-16T00:49:08-0400

"xa+yb=c"


"3x+2y=7"

"4x-y=2"


"x=1,y=2"

So, "c=a+2b"